a chemist mixes a 10% saline solution with a 20% saline solution to make 500 mililiters of a 16% solution. How many mililiters of each solution does the chemist mix together?
Answer by mananth(16949)

percent ---------------- quantity
Saline type I 10 ---------------- x ml
Saline type II 20 ------ 500 - x ml
Mixture 16.00% ---------------- 500

10 x + 20 ( 500 - x ) = 16 * 500

10x+10000-20x= 8000
10x-20x =8000-10000
-10x= -2000
/ -10
x=200 ml 10.00% Saline type I
300 ml 20.00% Saline type II

m.ananth@hotmail.ca